OpenStudy (anonymous):

I'm confused as to how to start this problem. Solve the di fferential equation 6xy dx + (4y + 9x^2)dy = 0 by fi nding an appropriate integrating factor to make it exact.

OpenStudy (anonymous):

I'd first change it so that's its not implicit..

OpenStudy (anonymous):

OK, so divide everything by dx, right?

OpenStudy (anonymous):

Well, technically speaking, you're not "dividing", but yeah..

OpenStudy (anonymous):

Actually there is probably an easier way.

OpenStudy (anonymous):

hold on let me confirm

OpenStudy (tkhunny):

What do you think? \(\dfrac{\partial}{\partial{y}}((6xy)\cdot r(x,y)) = \dfrac{\partial}{\partial{x}}((4y + 9x^{2})\cdot r(x,y))\) So, doing the drudgery: \([(6xy)\cdot r_{y}(x,y) + (6x)\cdot r(x,y)] = [(4y + 9x^{2})\cdot r_{x}(x,y) + (18x)\cdot r(x,y)]\) Now some algebra... Let's see what you get.

OpenStudy (anonymous):

Right, not really. OK, so if I do that, I get:\[6xy+(4y+9x^2)y'=0\]Then, I tested it for exactness by taking the partials of each term. I get \[6x\] and \[18x\] which are not equal. This fails the test for exactness, but I think they want me to use an integrating factor to change the terms into something that will be exact.

OpenStudy (anonymous):

OK, stand by, I'll churn the algebra wheel.

OpenStudy (anonymous):

OH! I see what you're saying! Give me a moment.

OpenStudy (anonymous):

To solve it, we can integrate both groups, compare the answers, and construct the solution from that: the integral of 6xy^3 dx is 3x^2y^3 + some function of y that would be obliterated by ordinary x differentiation. the integral of (4y^3+9x^2y^2)dy is y^4 + 3x^2y^3 + some function of x that would be obliterated by ordinary y differentiation. There's a term in common: the "hybrid" term 3x^2y^3 that has both x's and y's. The spare y^4 is part of the answer, too. I get y^4 + 3x^2y^3 = C where C is some constant as an implicitly defined solution.

OpenStudy (anonymous):

^ This was the easier way - found an explanation of it on yahoo

OpenStudy (anonymous):

OK, thanks king. I'm going to follow tkhunny's method through to the end, because it looks like the technique we discussed in class.

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

OK, tkhunny, I'm kind of getting lost in the sea of letters here, but it looks like isolating the integrating factor r(x,y) is what I should do next. Doing this, I get \[12xr()=6xyr _{y}()-(4y+9x^2)r_x()\] This looks like it could be a chain rule expansion, but I'm not sure. I could divide by the 12x, but until I know exactly where I'm going, I figured I'd leave it there with the r(). Am I on the right track?

OpenStudy (anonymous):

I have to step out for class, but I'll come back to this in a little while. Thank you all.